# Neighborhood Attention Transformer with Progressive Channel Fusion for Speaker Verification

### v1

@@ Abstract
Transformer-based architectures for speaker verification typically require more training data than ECAPA-TDNN. Therefore, recent work has generally been trained on VoxCeleb1&2. We propose a backbone network based on self-attention, which can achieve competitive results when trained on VoxCeleb2 alone. The network alternates between neighborhood attention and global attention to capture local and global features, then aggregates features of different hierarchical levels, and finally performs attentive statistical pooling. Additionally, we employ a progressive channel fusion strategy to expand the receptive field in the channel dimension as the network deepens. We trained the proposed PCF-NAT model on VoxCeleb2 and evaluated it on VoxCeleb1 and the validation sets of VoxSRC. The EER and minDCF of the shallow PCF-NAT are on average more than 20% lower than those of similarly sized ECAPA-TDNN. Deep PCF-NAT achieves an EER lower than 0.5% on VoxCeleb1-O.
@@ 1 Introduction
This paper introduces Neighborhood Attention [27], a concept from computer vision, into ASV. We alternate between neighborhood attention and global attention (original self attention) to capture local and global features. Adopting an architecture similar to ECAPA-TDNN, we concatenate the outputs of all blocks along the channel dimension and then apply attentive statistics pooling. Additionally, We use the progressive channel fusion strategy proposed by PCF-ECAPA-TDNN, replacing linear layers with grouped 1-dimensional convolutions with a kernel size of 1, and gradually reducing the number of groups to expand the receptive field in the channel dimension.
@@ 2.2 PCF-ECAPA-TDNN
PCF-ECAPA-TDNN uses group convolutions and gradually reduces the number of groups, thereby expanding the receptive field in the frequency (channel) dimension as the network deepens. This is known as the progressive channel fusion strategy. A key difference between ResNet and TDNN is that 2D convolutions in ResNet have a local receptive field in both the frequency and time dimensions, whereas TDNN maintains a global receptive field in the frequency dimension throughout, which increases the risk of overfitting due to the large number of parameters. Additionally, PCF-ECAPA-TDNN deepens the network by using four blocks, each containing two SE-Res2Blocks. Unlike the Res2Block in ECAPA-TDNN, PCF-ECAPA-TDNN adds a branch with a kernel size of 1.
@@ 3.2 Progressive Channel Fusion
To ensure that the receptive field in the channel dimension expands progressively with network depth, similar to the time dimension, and to reduce the number of parameters to mitigate the risk of overfitting, we use a large number of grouped 1D convolutions, similar to PCF-ECAPA-TDNN [3]. Specifically, we use four grouped 1D convolutions to process the input Fbank features, with the number of groups set to 8, 4, 2, and 1, respectively, and then feed the results into four blocks. The input to the second and subsequent blocks is augmented with the output from the previous block. For the four blocks, we replace the linear layers, including the $qkv$ layers, with grouped 1D convolutions, with the number of groups also set to 8, 4, 2, and 1. Unlike PCF-ECAPA-TDNN, we reduce the computational load by down-sampling the Fbank features in the time dimension by half, setting both the kernel size and stride of the convolution to 2.
@@ 4.2 Model Settings
We used the ECAPA-TDNN implemented by SpeechBrain [39] as the first baseline model, which includes three SE-Res2Blocks with 512 channels. These three blocks use dilated 1D convolutions with dilations of 2, 3, and 4, respectively.
@@
Following the original configuration, this model includes four blocks, each containing two SE-Res2Blocks with branches and 512 channels. The numbers of groups for the grouped 1D convolutions in the blocks are 8, 4, 2, and 1, respectively.
@@
Similar to PCF-ECAPA-TDNN, our proposed model consists of four blocks. The first and third blocks use neighborhood attention entirely, while the second and fourth blocks only use global attention in the last layer. Similar to MFA-Conformer, we use only four heads in the global attention layers to reduce memory usage and set the channel number to 256 for all blocks. For the neighborhood attention layers, the number of heads is set to a more efficient 16, with a window size of 27. Additionally, we scale the model by adjusting the number of attention layers in each block. For example, MFA-NAT (3x4) indicates that each block in the model contains three attention layers. Finally, the drop path rate for MFA-NAT models of different depths is set to 1.0.
@@
The proposed model includes four pairs of down-sampling layers and blocks, with the number of groups set to 8, 4, 2, and 1, respectively, to be consistent with PCF-ECAPA-TDNN. Unlike MFA-NAT, the two global attention layers of PCF-NAT are placed at the end of the first and third blocks. Additionally, PCF-NAT models of different depths use different drop path rates: PCF-NAT (3x4), (4x4), (5x4), and (6x4) use drop path rates of 1.0, 1.1, 1.2, and 1.3, respectively. For other configurations, PCF-NAT is identical to MFA-NAT.
@@ 4.4 Evaluation
Equal Error Rate (EER) and minimum Detection Cost Function (minDCF) were reported with CFA = CMiss = 1 for performance evaluation. The Ptarget of minDCF was set to 0.01 for VoxCeleb1 and 0.05 for VoxSRC. Furthermore, we calculate the number of batches processed per second by the models on a single Nvidia GPU 4090 to evaluate the inference speeds. Each batch contains 512 six-second utterances. We use PyTorch’s ‘torch.compile‘ with the default mode to accelerate model inference.
@@ 5.1 Results on VoxCeleb1 and VoxSRC
MFA-NAT (3X4) 12.6M 10.34 0.558 0.0731 0.738 0.0818 1.347 0.1328
MFA-NAT (4X4) 15.8M 8.32 0.580 0.0601 0.711 0.0777 1.305 0.1295
MFA-NAT (5X4) 18.9M 6.96 0.542 0.0792 0.721 0.0752 1.327 0.1351
MFA-NAT (6X4) 22.1M 5.98 0.601 0.0606 0.718 0.0749 1.305 0.1340
PCF-NAT (3X4) 7.6M 8.52 0.537 0.0500 0.703 0.0790 1.304 0.1332
PCF-NAT (4X4) 9.0M 6.82 0.526 0.0604 0.672 0.0750 1.260 0.1298
PCF-NAT (5X4) 10.5M 5.69 0.452 0.0392 0.672 0.0715 1.236 0.1270
PCF-NAT (6X4) 12.0M 4.87 0.436 0.0467 0.654 0.0712 1.198 0.1284
@@
MFA-NAT (3X4) 2.272 0.1214 2.671 0.1446 1.862 0.1260 3.447 0.1923
MFA-NAT (4X4) 2.229 0.1177 2.606 0.1453 1.815 0.1187 3.357 0.1845
MFA-NAT (5X4) 2.255 0.1190 2.634 0.1547 1.833 0.1241 3.334 0.1949
MFA-NAT (6X4) 2.232 0.1185 2.716 0.1588 1.764 0.1202 3.328 0.1852
PCF-NAT (3X4) 2.202 0.1209 2.609 0.1423 1.793 0.1262 3.340 0.1938
PCF-NAT (4X4) 2.163 0.1153 2.456 0.1403 1.723 0.1170 3.244 0.1830
PCF-NAT (5X4) 2.139 0.1120 2.723 0.1509 1.724 0.1164 3.307 0.1843
PCF-NAT (6X4) 2.059 0.1074 2.426 0.1405 1.723 0.1153 3.186 0.1830
@@
For scenarios requiring fast inference, the proposed MFA-NAT (3x4) is a good choice. Its throughput is 4.55% higher than ECAPA-TDNN, and its EER and minDCF are 6.86% lower on average than those of PCF-ECAPA-TDNN. By deepening the network, MFA-NAT (4x4) achieves EER and minDCF that are 3.57% lower on average than MFA-NAT (3x4), although its throughput is reduced by 19.54%. Further deepening of MFA-NAT does not yield better performance.
The use of the progressive channel fusion strategy allows PCF-NAT to achieve better performance and scalability. Compared to the EER and minDCF of MFA-NAT (4x4), PCF-NAT (4x4) shows an average reduction of 3.28%, while the deeper PCF-NAT (6x4) achieves an average reduction of 8.07%. However, in the current Pytorch implementation, grouped 1D convolutions are slower than linear layers despite having fewer parameters and computations. This results in less satisfactory inference speed for PCF-NAT, which uses many grouped 1D convolutions.
@@ 5.2 Ablation Study
Table 3: Ablation study of MFA-NAT (4x4). DCF is minDCF0.01. w/o is without. NA is neighborhood attention. GA is global attention.
@@
The effectiveness of the channel fusion strategy has been explained in the previous subsection. To understand the effects of other components used in the proposed model, We conducted an ablation study on MFA-NAT (4x3) and evaluated the results on VoxCeleb1, as shown in Table 3. In Experiment 2, we did not pad the sequence ends and restricted the neighborhood attention window to within the sequence range, which slightly increased the EER and minDCF. This supports our decision to follow the earlier Stand-Alone Self-Attention approach. In Experiment 3, where only neighborhood attention was used, the EER and minDCF increased by an average of 10.82%, indicating the usefulness of the two global attention layers in MFA-NAT. However, Experiment 4 showed that adding more global attention layers worsened the results, even if only the last layer of blocks 1 and 3 was changed to global attention. Experiment 5 replaced all batch normalization layers before multi-level feature aggregation with layer normalization, resulting in an average increase of 2.66% in EER and minDCF. In Experiment 7, removing MFA but retaining its subsequent dense layer led to a deterioration of 4.74% in the results.

### v2

@@ Abstract
Transformer-based architectures for speaker verification typically require more training data than ECAPA-TDNN. Therefore, recent work has generally been trained on VoxCeleb1&2. We propose a backbone network based on self-attention, which can achieve competitive results when trained on VoxCeleb2 alone. The network alternates between neighborhood attention and global attention to capture local and global features, then aggregates features of different hierarchical levels, and finally performs attentive statistics pooling. Additionally, we employ a progressive channel fusion strategy to expand the receptive field in the channel dimension as the network deepens. We trained the proposed PCF-NAT model on VoxCeleb2 and evaluated it on VoxCeleb1 and the validation sets of VoxSRC. The EER and minDCF of the shallow PCF-NAT are on average more than 20% lower than those of similarly sized ECAPA-TDNN. Deep PCF-NAT achieves an EER lower than 0.5% on VoxCeleb1-O. The code and models are publicly available at https://github.com/ChenNan1996/PCF-NAT.
@@ 1 Introduction
This paper introduces Neighborhood Attention [27], a concept from computer vision, into ASV. We alternate between neighborhood attention and global attention (original self-attention) to capture local and global features. Adopting an architecture similar to ECAPA-TDNN, we concatenate the outputs of all blocks along the channel dimension and then apply attentive statistics pooling. Additionally, We use the progressive channel fusion strategy proposed by PCF-ECAPA-TDNN, replacing linear layers with 1-dimensional group convolutions with a kernel size of 1, and gradually reducing the number of groups to expand the receptive field in the channel dimension.
@@ 2.2 PCF-ECAPA-TDNN
PCF-ECAPA-TDNN uses group 1D convolutions and gradually reduces the number of groups, thereby expanding the receptive field in the frequency (channel) dimension as the network deepens. This is known as the progressive channel fusion strategy. A key difference between ResNet and TDNN is that 2D convolutions in ResNet have a local receptive field in both the frequency and time dimensions, whereas TDNN maintains a global receptive field in the frequency dimension throughout, which increases the risk of overfitting due to the large number of parameters. Additionally, PCF-ECAPA-TDNN deepens the network by using four blocks, each containing two SE-Res2Blocks. Unlike the Res2Block in ECAPA-TDNN, PCF-ECAPA-TDNN adds a branch with a kernel size of 1.
@@ 3.2 Progressive Channel Fusion
To ensure that the receptive field in the channel dimension expands progressively with network depth, similar to the time dimension, and to reduce the number of parameters to mitigate the risk of overfitting, we use a large number of 1D group convolutions, similar to PCF-ECAPA-TDNN [3]. Specifically, we use four 1D group convolutions to process the input Fbank features, with the number of groups set to 8, 4, 2, and 1, respectively, and then feed the results into four blocks. The input to the second and subsequent blocks is augmented with the output from the previous block. For the four blocks, we replace the linear layers, including the $qkv$ layers, with 1D group convolutions, with the number of groups also set to 8, 4, 2, and 1. Unlike PCF-ECAPA-TDNN, we reduce the computational load by down-sampling the Fbank features in the time dimension by half, setting both the kernel size and stride of the convolution to 2.
@@ 4.2 Model Settings
We used the ECAPA-TDNN implemented by SpeechBrain [39] as the first baseline model, which includes three SE-Res2Blocks with 512 channels. These three blocks use 1D dilated convolutions with dilations of 2, 3, and 4, respectively.
@@
Following the original configuration, this model includes four blocks, each containing two SE-Res2Blocks with branches and 512 channels. The numbers of groups for the 1D group convolutions in the blocks are 8, 4, 2, and 1, respectively.
@@
Similar to PCF-ECAPA-TDNN, our proposed model consists of four blocks. The first and third blocks use neighborhood attention entirely, while the second and fourth blocks only use global attention in the last layer. Similar to MFA-Conformer, we use only four heads in the global attention layers to reduce memory usage and set the channel number to 256 for all blocks. For the neighborhood attention layers, the number of heads is set to a more efficient 16, with a window size of 27. Additionally, we scale the model by adjusting the number of attention layers in each block. For example, MFA-NAT (3$\times$4) indicates that each block in the model contains three attention layers. Finally, the drop path rate for MFA-NAT models of different depths is set to 1.0.
@@
The proposed model includes four pairs of down-sampling layers and blocks, with the number of groups set to 8, 4, 2, and 1, respectively, to be consistent with PCF-ECAPA-TDNN. Unlike MFA-NAT, the two global attention layers of PCF-NAT are placed at the end of the first and third blocks. Additionally, PCF-NAT models of different depths use different drop path rates: PCF-NAT (3$\times$4), (4$\times$4), (5$\times$4), and (6$\times$4) use drop path rates of 1.0, 1.1, 1.2, and 1.3, respectively. For other configurations, PCF-NAT is identical to MFA-NAT.
@@ 4.4 Evaluation
Equal Error Rate (EER) and minimum Detection Cost Function (minDCF) were reported with CFA = CMiss = 1 for performance evaluation. The Ptarget of minDCF was set to 0.01 for VoxCeleb1 and 0.05 for VoxSRC. Furthermore, we calculate the number of batches processed per second by the models on a single Nvidia GPU 4090 to evaluate the inference speeds. Each batch contains 512 six-second utterances. We use PyTorch’s $torch.compile$ with the default mode to accelerate model inference.
@@ 5.1 Results on VoxCeleb1 and VoxSRC
MFA-NAT (3$\times$4) 12.6M 10.34 0.558 0.0731 0.738 0.0818 1.347 0.1328
MFA-NAT (4$\times$4) 15.8M 8.32 0.580 0.0601 0.711 0.0777 1.305 0.1295
MFA-NAT (5$\times$4) 18.9M 6.96 0.542 0.0792 0.721 0.0752 1.327 0.1351
MFA-NAT (6$\times$4) 22.1M 5.98 0.601 0.0606 0.718 0.0749 1.305 0.1340
PCF-NAT (3$\times$4) 7.6M 8.52 0.537 0.0500 0.703 0.0790 1.304 0.1332
PCF-NAT (4$\times$4) 9.0M 6.82 0.526 0.0604 0.672 0.0750 1.260 0.1298
PCF-NAT (5$\times$4) 10.5M 5.69 0.452 0.0392 0.672 0.0715 1.236 0.1270
PCF-NAT (6$\times$4) 12.0M 4.87 0.436 0.0467 0.654 0.0712 1.198 0.1284
@@
MFA-NAT (3$\times$4) 2.272 0.1214 2.671 0.1446 1.862 0.1260 3.447 0.1923
MFA-NAT (4$\times$4) 2.229 0.1177 2.606 0.1453 1.815 0.1187 3.357 0.1845
MFA-NAT (5$\times$4) 2.255 0.1190 2.634 0.1547 1.833 0.1241 3.334 0.1949
MFA-NAT (6$\times$4) 2.232 0.1185 2.716 0.1588 1.764 0.1202 3.328 0.1852
PCF-NAT (3$\times$4) 2.202 0.1209 2.609 0.1423 1.793 0.1262 3.340 0.1938
PCF-NAT (4$\times$4) 2.163 0.1153 2.456 0.1403 1.723 0.1170 3.244 0.1830
PCF-NAT (5$\times$4) 2.139 0.1120 2.723 0.1509 1.724 0.1164 3.307 0.1843
PCF-NAT (6$\times$4) 2.059 0.1074 2.426 0.1405 1.723 0.1153 3.186 0.1830
@@
For scenarios requiring fast inference, the proposed MFA-NAT (3$\times$4) is a good choice. Its throughput is 4.55% higher than ECAPA-TDNN, and its EER and minDCF are 6.86% lower on average than those of PCF-ECAPA-TDNN. By deepening the network, MFA-NAT (4$\times$4) achieves EER and minDCF that are 3.57% lower on average than MFA-NAT (3$\times$4), although its throughput is reduced by 19.54%. Further deepening of MFA-NAT does not yield better performance.
The use of the progressive channel fusion strategy allows PCF-NAT to achieve better performance and scalability. Compared to the EER and minDCF of MFA-NAT (4$\times$4), PCF-NAT (4$\times$4) shows an average reduction of 3.28%, while the deeper PCF-NAT (6$\times$4) achieves an average reduction of 8.07%. However, in the current Pytorch implementation, 1D group convolutions are slower than linear layers despite having fewer parameters and computations. This results in less satisfactory inference speed for PCF-NAT, which uses many 1D group convolutions.
@@ 5.2 Ablation Study
Table 3: Ablation study of MFA-NAT (4$\times$4). DCF is minDCF0.01. w/o is without. NA is neighborhood attention. GA is global attention.
@@
The effectiveness of the channel fusion strategy has been explained in the previous subsection. To understand the effects of other components used in the proposed model, We conducted an ablation study on MFA-NAT (4$\times$4) and evaluated the results on VoxCeleb1, as shown in Table 3. In Experiment 2, we did not pad the sequence ends and restricted the neighborhood attention window to within the sequence range, which slightly increased the EER and minDCF. This supports our decision to follow the earlier Stand-Alone Self-Attention approach. In Experiment 3, where only neighborhood attention was used, the EER and minDCF increased by an average of 10.82%, indicating the usefulness of the two global attention layers in MFA-NAT. However, Experiment 4 showed that adding more global attention layers worsened the results, even if only the last layer of blocks 1 and 3 was changed to global attention. Experiment 5 replaced all batch normalization layers before multi-level feature aggregation with layer normalization, resulting in an average increase of 2.66% in EER and minDCF. In Experiment 7, removing MFA but retaining its subsequent dense layer led to a deterioration of 4.74% in the results.